Log4x2a=3 How do I solve this?


1
Expert's answer
2021-02-16T12:40:24-0500

Solution.

"log_{4x}{2^a}=3"

Find the domain of the function:

"4x>0, 4x\neq 1\implies \newline x>0,x\neq \frac{1}{4}," and

"2^a>0\implies a \in R."

Solve the equation:


"(4x)^3=2^a,""x^3=\frac{2^a}{4^3},""x^3=2^{a-6},""x=\sqrt[3]{2^{a-6}}."

"\sqrt[3]{2^{a-6}} \neq \frac{1}{4},"

"\frac{a-6}{3}\neq -2,"

"a\neq 0."

Answer. "x=\sqrt[3]{2^{a-6}}, \text{if } a\neq 0."